# Beyond the Carnot limit: work extraction via an entropy battery

### v1

@@ Abstract
We explore the consequences of generalized thermodynamics, as interpreted from the perspective of information theory, on an ensemble’s capacity to extract energetic work. We demonstrate that by utilizing unitary heat engines of different kinds, we can capitalize on the full entropy capacity of the ensemble across multiple conserved quantities. This leads to the extraction of energetic work with efficiencies surpassing the standard Carnot efficiency limit, all the while running at maximum power. This is achieved without requiring modified thermal sources, which is a method commonly used in the field of quantum heat engines. We pay particular attention to spin angular momentum, and show that when treated as a mutually independent conserved quantity, spin baths exhibit thermodynamic behaviour analogous to conventional heat baths, with corresponding quantities and models. This includes spin-analogous heat capacities, Einstein solid and Debye models, entropic responses, and Bose-Einstein and Fermi-Dirac statistics. Spin baths also follow the fluctuation-dissipation theorem. Further, we examine the role of particle statistics in determining the maximum entropy capacity of the battery. In doing so, we argue that indistinguishability is necessary for interactions to occur, motivating a perspective in which information itself is treated as the indistinguishable property of particles, with information (in the form of coherence) transferable between both physical degrees of freedom such as energy and spin, and particles when the system acts unitarily. This work will find application in fields that require near energy degenerate spin statistics, such as in spinor Bose–Einstein condensates and spintronics. Further, our method will have implications for the fields of quantum heat engines, quantum batteries, quantum error correction, and information and resource theories.
@@ I Introduction
Born out of the Industrial Revolution, heat engines provided not only practical utility but, more importantly, a theoretical framework for describing some of the most fundamental principles of nature. The pursuit of increasingly powerful and efficient engines, refrigerators, and related devices, driven by both technological ambition and the intellectual curiosity of figures such as Carnot and Clausius [1], led to the formulation of the four laws of thermodynamics, which govern the dynamics of heat and energy. The need to explain the emergence of these laws from the microscopic constituents of thermodynamic systems lead to statistical mechanics, a theory that successfully described the macroscopic properties of the ensemble using probability and statistical techniques. Maxwell, by utilizing statistical mechanics, proposed a mechanism by which an entity with sufficient knowledge of the microscopic states could violate the laws of thermodynamics [2], in what came to be known as Maxwell’s demon.
@@
Szilard [3], Landauer [4], and Bennett [5] each contributed, at times indirectly, to the resolution of this paradox, with Bennett ultimately recognising that such a demon would only operate for as long as it’s memory has capacity to store information about the state of the system. If this capacity were saturated, as would occur for a demon constrained by finite physical resources, no further work could be extracted from the system. Bennett showed, using Landauer’s principal, that the energetic cost of erasing the demon’s memory would be greater-than or equal-to the amount of work extracted during the engine cycles, therefore closing the paradox.
@@
The roles of Information and entropy (the latter being the lack of the former) in thermodynamics became a recurring theme in the study of thermodynamic systems, even prior to the resolution of the demon paradox. Jaynes, from the perspective of information theory, generalized statistical mechanics to include arbitrary conserved quantities, therefore removing the special treatment given to energy [6, 7]. His rational can be intuitively understood by recognising that the time invariant nature of conserved quantities must allow for the storage of information. Vaccaro and Barnett later used Jaynes’s principle to show how the erasure of information in the Maxwell demon thought experiment can be performed using spin angular momentum [8, 9], leading to the development of memory powered heat engines [10] and gas spin heat engines [11]. The application of Jayne’s work to quantum systems has also led to developments in generalized heat baths, quantum batteries, and resource theories [12, 13, 14]. With a renewed focus on information as a general resource, it appears as though thermodynamics is not describing the flow of energy and heat, but instead the flow of information and entropy at the most fundamental level, with conserved quantities acting merely as the physical degrees of freedom for which the information is encoded. Yet, the concept of generalized information as a resource has not been widely utilized by the scientific community, as no clear practicality has emerged for applying thermodynamics in this manner. This paper provides one such practicality.
To be precise, the term “information” is coherence, corresponding to certainty in the system’s state. For instance, a pure state conveys unambiguous information about a particle’s position in state space and thus the coherence of the state is a manifestation of information. For an ensemble of interacting particles, information is extended to include coherent correlations, or entanglement. Entropy on the other hand is a measure of the “mixed”, or classical uncertainty of the conserved quantity, with entropy representing the lack of information about the particle’s state. Whether this entropy is an ‘anthropomorphic’ by-product of our ignorance as an observer [15], or is as physical as the conserved quantities themselves, the act of obtaining information and thus reducing the systems entropy takes physical resources [5, 8], thus entropy must be treated as a resource itself.
The promises of modern resource theories, quantum batteries, and quantum heat engines include: collective power enhancement in the charging of quantum batteries [16], heat engines utilizing non-energetic cold baths [10, 11], the saturation and surpassing of the classical Carnot energy efficiency limit [17, 18], and methods of error correction in quantum computers [19, 20]. Information and resource theories have also found broad application across the fields of quantum state amplification and metrology [21, 22, 23]. Of the works presented thus far, none have explicitly explored how the entropy capacity of an ensemble changes once other conserved quantities are introduced. More importantly, none have explored how coherent coupling mechanisms between conserved quantities, such as those used by [10, 11], change both the efficiency of energetic work extraction and the total work extractable. [24, 25, 12] have independently explored the reversible exchange of conserved quantities within the framework of generalized thermodynamics; however, they did not address the extent to which these methods influence the practicality of applying such principles to large ensembles for performing useful work.
Furthermore, a common approach used in the field of quantum heat engines to surpass the Carnot efficiency limit is to exploit intrinsic quantum coherence in the thermal bath, for example via squeezed-thermal sources [26, 27, 28], or through induced quantum coherence [29, 30, 31]. However, introducing coherence into a thermal system incurs an energy cost under the second law of thermodynamics, as such non-thermal states are ‘purer’ than the corresponding thermal state at the same temperature, which by definition maximizes entropy. Taking the totality of the resources required to run such quantum heat engines above the Carnot limit, one finds no such energy efficiency gain was made. This is reflected in the fact that such devices operate below the generalized Carnot limit [32], which accounts for the efficiency gains achievable with engineered non-thermal states. Furthermore, the field of quantum heat engines is often limited in total extractable work as it generally relies on individual quantum systems, such as single ions or molecules, where quantum effects significantly influence the engine cycle. In general, the Carnot efficiency limit for two thermal sources has prevailed as unsurpassable, with quantum systems only possibly saturating the limit [17].
We argue, in the spirit of Jaynes, Vaccaro and others, that a reinterpretation of thermodynamics into a purely informational basis allows for an intuitive interpretation of the operation of all adiabatic and unitary heat engines, with them now acting as mechanisms that transfer coherent information and incoherent entropy across every conserved quantity. This interpretation applies also to the engines we use in what we call the entropy battery. The entropy battery, and this paper, will also consider conserved quantities that meet the criteria of being experimentally accessible and controllable.
@@
There are 8 conserved quantities for which one can in principle encode information. This is because there exist eight fundamental physical symmetries, each associated with a corresponding conservation law and conserved quantity, in accordance with Noether’s theorem [33, 34]. The known conserved quantities are presented in Table 1.
@@
Table 1: The known conserved quantities as derived from continuous physical symmetries according to Noether’s theorem, excluding Baryon, Lepton difference. Also shown are the known analogous definitions of ‘heat’ and ‘work’ for these conserved quantities, with $\mathbb{Q}$ and $\mathbb{W}$ in place for the quantities without literature definitions. Energy, linear momentum, and the Lorentz 3-vector boost may be defined on both discrete and continuous state spaces, spin angular momentum, electric charge, color charge, and weak isospin are intrinsically discrete and bounded. The generalized heat and work are assumed to exist for these conserved quantities based on the commentary in this paper, although the application of these quantities as useful thermodynamic properties is questionable with the current state of technology.
Among the eight conserved quantities, only energy and spin angular momentum satisfy the criteria of accessibility and controllability, with systems such as Bose–Einstein condensates and trapped atoms allowing controllable preparation of both thermal motion and internal spin states. We will therefore focus our efforts in this paper on energy in a single degree of freedom and the z-component of spin angular momentum, without loss of generality.
We anticipate that the entropy battery will have applications extending beyond future high-efficiency heat engines and batteries. For example, heat storage in environments where waste heat removal is challenging (e.g., the vacuum of space), mitigation of entropy-induced noise in quantum measurements such as those in quantum computing, and refrigeration with efficiencies surpassing those of present technologies. Separately, our analysis of information thermodynamics applied to spin angular momentum will have application in fields that require a description of ensemble spin state statistics in energy degenerate regimes, such as spinor BEC’s and spintronics.
The results of the entropy battery and the corresponding spin thermodynamics are presented in §II, followed by a discussion of entropy capacity maximization in §III. Model derivations are provided in the Appendix.
II The entropy battery
In Fig.1 we present an overview of the entropy transfer process during the operation of the entropy battery. In our model, we assume the energy and z-component of spin angular momentum conserved quantities are mutually independent, making the entropy of the ensemble initially extensive over energy $\mathbb{S}_{\text{energy}}$ and spin $\mathbb{S}_{\text{spin}}$,
Since spin angular momentum is independent of any other quantity in our model, it behaves as an independent thermodynamic system. Table 2 summarizes the analogous spin-angular-momentum thermodynamic relations derived in this paper, with derivations provided in Appendices AE, including the four laws of generalized thermodynamics.
@@ II The entropy battery
Bosonic heat for solids in
the thermodynamic limit
@@
Table 2: Summary of the spin-analogous thermodynamic properties. In these expressions, $\braket{\bar{J}_{z}}=\braket{J_{z}}/\hbar$ is the unitless average angular momentum, $d_{m}$ is number of non-degenerate states for the energy, spin $m=E,s$ degrees of freedom respectively, NN is the number of particles, $\varepsilon_{j}=h\nu_{j}$ is the energy of the $j$th vibrational mode with frequency $\nu_{j}$, and $\varepsilon_{D}=h\nu_{D}$ is the Debye energy. Where relevant we have also included references. The waste response is a generalization of the heat capacity to arbitrary conserved quantities, and is coined based on the results of the fluctuation-dissipation theorem, whereby $C_{V}$ and $C_{s}$ are response functions (2). The derivation for each of these expressions for the spin angular momentum conserved quantity is presented in appendices A-D. The first and second laws of generalized thermodynamics apply to closed systems, which is the scenario used in the entropy battery model. The third law describes the generalized thermalization condition of two canonical ensembles.
Figure 1: Overview of (a) a conventional Carnot heat engine between hot and cold baths, and (b) the entropy battery. In (b), the equivalent conserved quantity ‘heat’ baths with general conserved quantity $\braket{A_{m}}$ are represented by black cubes aligned vertically. Energy couples to spin angular momentum within the entropy battery via a unitary “heat engine,” realized experimentally through coherent Raman transitions [10, 11]. Spin labor $\mathcal{L}$ supplied by the entropy battery enables the conversion of heat energy into work. There aren’t any equivalent coupling mechanisms for the other conserved quantities. Entropy from the environmental heat bath is transferred into the battery’s internal baths (energy, spin, etc.), enhancing work extraction until their entropy capacities are saturated or equilibrium is reached. As these other conserved quantity ‘heat’ baths reside within the battery’s particles, the scheme achieves a higher volumetric work density than conventional heat engines.
In the entropy battery model (Fig.II), we assume the $N_{\text{res}}$ particles that make up the battery have $d_{s}$ energy degenerate internal spin states and $d_{E}$ evenly spaced external (harmonic position, momentum) energy modes, and as such the ensemble has an entropy capacity determined by equations (2), (2) and (2) for distinguishable, bosonic or fermionic particles respectively. Energy and spin angular momentum are only coupled in the presence of a magnetic field $B_{z}$, as such we assume $|B_{z}|=0$. Even if this is only approximately true, our analysis will continue to apply for as long as the Zeeman splitting $|\delta_{B}|=g\mu_{0}B_{z}$ is much less than the harmonic spacing $h\nu$ as to avoid inadvertent coupling of the spin and energy degrees of freedom. Each particle in the battery has spin $S\hbar$, where $d_{s}=2S+1$, and spin state eigenvalues $m_{j}=(j-S)\hbar$, or equivalently labeled in the computational basis $j$ as is often done in this paper, where $j\in\{0,1,\dots,2S\}$. The distribution of particles across the $j$ spin states is governed by Bose-Einstein statistics for bosons and Fermi-Dirac statistics for fermions, depending on their overall wavefunction symmetry upon exchange, as derived in appendices B.2, B.3. Table 2 also summarises these distributions, along with the generalized heat capacities (waste responses), and the entropy capacities for both energy and spin angular momentum.
What is notable upon viewing Table 2 is that an ensemble with only a spin angular momentum degree of freedom (energy degenerate system) still contains thermodynamic properties. If we pay particular attention to the heat capacity equations, (2), (2)-(2), then it would appear as though spin angular momentum can store heat similarly to an energy heat bath, or rather the spin-analogous of heat. This is where we expand the concept of heat to generally ‘waste’, which we define as incoherence within an arbitrary conserved quantity. According to the fluctuation-dissipation theorem [36] and (2), $C_{s}$ and $C_{V}$ are also response functions of the ensemble. For this reason, we introduce the term “waste response” as a generalized replacement for heat capacity across arbitrary conserved quantities. A similar argument motivates our naming of the entropic response in eq.(2). When independent of all other properties, the spin degree of freedom of the particles collectively forms an effective “waste” bath, separate from the energy heat bath. This spin waste bath does not store heat energy directly; rather, as noted above, it stores the spin equivalent of heat, the spin therm $\mathcal{Q}$. One might wonder then if the entropy associated with the spin waste bath immediately affects an ensembles ability to extract energetic work. When in the presence of a magnetic field that couples the internal spin to the external energy modes, one typically does observe an impact, with this documented as the Schottky anomaly [38]. However, as there is an energy cost in generating the magnetic field, then there is no advantage in work extraction. We also find no advantage when energy and spin angular momentum degrees of freedom are mutually independent, $|B_{z}|\rightarrow 0$, as spin/energy collisions cannot exchange heat between the two conserved quantities, by definition of mutual independence. This is despite the real extensive entropy of the spin-energy system when mutually independent (1). It is only upon introducing the unitary “heat engine” that a genuine work advantage arises.
@@
Before proceeding, we note that the Raman-based coupling heat engine incurs a non-energetic resource cost. Specifically, the engine requires coherence to convert heat into work, as dictated by the second law of generalized thermodynamics (2). In our model, this coherence is provided by spin labor, $\mathcal{L}$, which corresponds physically to the polarization of the spin particles within the battery reservoir. Thus, the entropy battery can be viewed as a device that exploits general quantum resources, in this case coherent spin angular momentum, to enhance work extraction efficiency.
II.1 Specifics of the entropy battery

### v2

@@ Abstract
Heat is a physical manifestation of entropy, where the removal of entropy from a thermal energy reservoir permits the conversion of heat into work. This entropy transfer is facilitated by the cold thermal energy reservoir in typical heat engines. Recent developments in quantum heat engines that operate between thermal energy and spin angular momentum reservoirs show that it is possible to transfer entropy out of energy and into a different conserved quantity. The implications of this type of entropy transfer have not been fully explored, especially on the work extractable using an ensemble with multiple conserved quantities. Using the aforementioned heat engines, we show that such an ensemble will transform heat into work beyond the Carnot efficiency limit while operating at maximum power. This result is obtained without induced quantum coherence, a technique commonly used in the field of quantum heat engines to achieve the same outcome. Without loss of generality, we also show that thermal spin reservoirs behave as thermodynamic baths with well-defined temperatures, heat capacities, and fluctuation–dissipation relations. Finally, our analysis of entropy capacity suggests that particle indistinguishability is necessary for inter-particle interactions, and for entropy to transfer between canonical ensembles. These results establish a foundation for entropy-based quantum devices that can extract work from a hot thermal energy reservoir more efficiently than possible with a cold thermal energy reservoir. These devices also act as high energy-density batteries and efficient heat storage systems. Our results will have implications for quantum heat engines, spinor condensates, spintronics, and quantum batteries.
@@ I Introduction
Heat engines provide not only practical utility, but also a theoretical framework for describing some of the most fundamental principles of nature. The pursuit of increasingly powerful and efficient engines, refrigerators, and related devices, driven by both technological ambition and the intellectual curiosity of figures such as Carnot and Clausius [1], led to the formulation of the four laws of thermodynamics. The need to explain the emergence of these laws from the microscopic constituents of thermodynamic systems lead to statistical mechanics, a theory that successfully described the macroscopic properties of the ensemble using probability and statistical techniques. Maxwell, by utilizing statistical mechanics, proposed a mechanism by which an entity with sufficient knowledge of the microscopic states could violate the second law [2], in what came to be known as Maxwell’s demon.
@@
Szilard [3], Landauer [4], and Bennett [5] each contributed, at times indirectly, to the resolution of the demon paradox, with Bennett ultimately recognising that such a demon would only operate for as long as it’s memory has capacity to store information about the state of the system. If this memory capacity were saturated, as would occur for a demon constrained by finite physical resources, no further work could be extracted from the system. Bennett showed, using Landauer’s principal, that the energetic cost of erasing the demon’s memory would be greater-than or equal-to the amount of work extracted during the engine cycles, therefore closing the paradox.
@@
The roles of Information and entropy (the latter being the lack of the former) in thermodynamics became a recurring theme in the study of thermodynamic systems, even prior to the resolution of the demon paradox. Jaynes, from the perspective of information theory, generalized statistical mechanics to include arbitrary conserved quantities, therefore removing the special treatment given to energy [6, 7]. His rational can be intuitively understood by recognizing that the time invariant nature of conserved quantities must allow for the encoding of information. Vaccaro and Barnett later used Jaynes’s principle to show how the erasure of information in the Maxwell demon thought experiment can be performed using spin angular momentum [8, 9, 10, 11], leading to the development of quantum spin heat engines [12, 13] and gas spin heat engines [14]. The application of Jaynes’s work to quantum systems has also led to developments in generalized heat baths, quantum batteries, and resource theories [15, 16, 17]. It would now appear as though thermodynamics is not describing energy and heat, but instead information and entropy. However, unlike energy and heat, information and entropy can transfer between conserved quantities according to [8, 9, 10, 11, 12, 13, 14, 15, 16], where the conserved quantities act as the physical degrees of freedom for which information is encoded. The concept of freely transferrable entropy across conserved quantities has not been widely utilized by the scientific community as no clear practicality has emerged for applying thermodynamics in this manner. This article provides one such practicality.
This article is structured in the following manner. In §II we discuss our rational for treating entropy as a conserved quantity in isolated systems. In §III we discuss the current applications of quantum information and the open question we are investigating. In §IV we introduce the entropy battery, followed by §V we present the energy efficiency of the battery. In this section we also summarize the thermodynamic properties of spin angular momentum, which are derived in the Appendix. In §VI we discuss the implications of this type of device, and finally, in §VII we present methods of maximizing the batteries entropy capacity.
II Entropy as a conserved quantity
In quantum mechanics, isolated systems evolve unitarily and the system’s total entropy is conserved $\delta\mathbb{S}=0$. Subsystem interactions do not produce entropy, even when those subsystems are at non-equilibrium. This fact is formalized in the ‘no-hiding’ [18], ‘no-cloning’ [19], and ‘no-deletion’ [20] theorems, all of which point to the impossibility of creating or destroying quantum information, with early experimental evidence [21, 22] agreeing. Despite the evidence of the conservation of information and entropy, there is often misconceptions of whether entropy production can occur in isolated systems. This usually stems from the original formulation of the second law $\delta\mathbb{S}\geq 0$, and reinforced by Boltzmann’s statistics and associated entropy [23]. However, as already discussed, no such entropy production can occur if the system is isolated. Any perception of entropy production in quantum systems is due to two common misconceptions,
Misconception 1.
A system’s total entropy after subsystem interaction can be calculated as a sum of the subsystem entropies.
Misconception 2.
An isolated system can decohere, and therefore a systems entropy can increase.
Misconception 1 is usually made if subsystems are assumed to be independent at all times. If the subsystems interact, with interactions necessarily correlating (entangling) the subsystems, then this assumption is clearly false. Entanglement leads to the inseparability of subsystem wavefunctions and entropy sub-additivity $\sum_{i}\delta\mathbb{S}_{i}\geq\delta\mathbb{S}=0$, where $\mathbb{S}_{i}=-\text{Tr}[\rho_{i}\ln(\rho_{i})]$ is the entropy of subsystem $i$ with reduced density matrix $\rho_{i}=\text{Tr}_{j}[\rho]$, and $\rho$ is the system density matrix. The sub-additivity of subsystem entropies creates the appearance of entropy production in the total system. Without misconception 1, we return to the true change in system entropy $\delta\mathbb{S}=0$.
Misconception 2, although less common than 1, comes from a misunderstanding as to the origin of decoherence. Decoherence only occurs when a system interacts with an external environment, and it has been shown both theoretically [24] and to some extent experimentally [21] that unitary evolution of system+environment creates the appearance of decoherence in the system. With decoherence typically leaving the system in a thermal state after a sufficiently long time, one often incorrectly concludes that entropy production can occur in isolated, decohering systems.
There are also cases in classical mechanics where entropy production causes inconsistencies. The Gibbs paradox, although seemingly unrelated, points to the issue of entropy production for two ideal gases that are arbitrarily similar. This problem is only resolved once particle indistinguishability is accepted [25]. Particle indistinguishability relates directly to the permutation of information (a parity operation), or equally to the transferability of information between particles. The resolution to the Gibbs paradox, along with the requirement of photon indistinguishability to resolve the UV catastrophe, the indistinguishability of Bosons and Fermions in statistical mechanics, the behavior of heat capacities at low temperature, and inherent indistinguishability in quantum mechanics suggests that entropy production in macroscopic systems is also illusionary. Instead, for all of the evidence to be consistent, then entropy and information must be conserved, with them only able to transfer between the subsystem of an isolated system. Whether the true nature entropy of as an ‘anthropomorphic’ by-product of our ignorance as an observer [26], or is as physical as the conserved quantities themselves, the act of obtaining information and reducing the systems entropy takes physical resources [27, 9, 8], then entropy must be treated as a resource itself.
For these reasons, we treat information and entropy as conserved quantities, with entropy and information manifesting in thermodynamic systems as generalized heats and works. However, while physical conserved quantities such as energy are constrained to their domains, entropy and information can in principle reversibly traverse the physical conserved quantities [9, 8, 15, 12, 13, 14], meaning the generalized heats and works in each of the conserved quantities can reversibly exchange. We will discuss this further in §IV. This idea points to an intriguing possibility: the exchange of entropy from a conserved quantity we wish to utilize, such as energy, into another conserved quantity that can be discarded, such as spin.
III The application of quantum information
The fields of modern resource theories, quantum batteries, and quantum heat engines all make use of quantum information to varying extent. They provide pathways towards quantum advantage in quantum thermodynamic systems. These advantages include: collective power enhancement in the charging of quantum batteries [28], heat engines utilizing non-energetic cold baths [13, 12, 14], non-classical and non-energetic methods of information erasure [9, 8], the saturation and surpassing of the classical Carnot energy efficiency limit [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39], and methods of error correction in qubits [40, 41]. Information and resource theories have separately found broad application in the fields of quantum state amplification and metrology [42, 43, 44]. This list of applications, although hardly exhaustive, does point to some of the main ways quantum information principles are used in real systems.
It should be noted that the observation of super-Carnot efficiencies in certain quantum heat engines has only been achieved by using modified (squeezed, induced coherence) thermal energy reservoirs, with all such engines operating below the generalised Carnot limit [45]. These non-thermal states require resources (typically energy) to prepare. This initial resource cost reduces the total efficiency of the heat engine below the Carnot limit. It is our assessment that, except for heat engines that utilize multiple conserved quantities (spin angular momentum and energy in the case of [13, 12]), no quantum heat engine has truly surpassed the Carnot limit.
As for the transfer of entropy between conserved quantities, [9, 11, 10, 15] have independently explored the reversible exchange of conserved quantity entropies within the framework of generalized thermodynamics. Meanwhile, [12, 13, 14] provide explicit unitary operations that couple spin angular momentum and energy, operations that we later use in §IV for the entropy battery. These studies have primarily focused on the principles of entropy transfer or the actual heat engines that run between two conserved quantities (energy and spin) on single quantum systems (quantum dots or trapped ions). Therefore, to the best of our knowledge, there has been very little exploration of an ensemble’s capacity to efficiently store entropy over multiple conserved quantities, and the effect this has on the efficiency of energetic work extraction.
@@ III The application of quantum information
There are eight conserved quantities for which information can in principle be encoded, one for each physical symmetry according to Noether’s theorem [46, 47]. The known conserved quantities are presented in Table 1.
@@
Table 1: The known conserved quantities as derived from continuous physical symmetries according to Noether’s theorem, excluding Baryon, Lepton difference. Also shown are the known analogous definitions of ‘heat’ and ‘work’ for these conserved quantities, with $\mathbb{Q}$ and $\mathbb{W}$ in place for the quantities without literature definitions. Energy, linear momentum, and the Lorentz 3-vector boost may be defined on both discrete and continuous state spaces, while spin angular momentum, electric charge, color charge, and weak isospin are intrinsically discrete and bounded. The generalized heat and work are assumed to exist for these conserved quantities based on the commentary in this paper, although the application of these quantities as useful thermodynamic properties is questionable with the current state of technology.
Of the eight conserved quantities, only energy and spin angular momentum are accessible and controllable by bench top experiments, with systems such as Bose–Einstein condensates and individually trapped ions allowing controllable preparation of both thermal motion and internal spin states. We will therefore focus our efforts in this article on energy in a single motional axis and the z-component of spin angular momentum, with our analysis applying equally to any conserved quantity.
IV The entropy battery
The entropy battery is a device that capitalizes on entropy conservation, although physically it is no more than a canonical ensemble with multiple conserved quantities. We assume the container for which the particles exist in is a quadratic potential, such that each particle acts as a quantum harmonic oscillator with energy spacing $h\nu$. We then suppose each particle has internal, energy degenerate spin states. These spin states correspond to the quantized values of the $z$-component of spin angular momentum. The canonical ensemble as described therefore contains two conserved quantities, which can be made mutually independent by setting the Zeeman splitting to 0, $\delta_{B}=0$. Our analysis will continue to apply even for $|\delta_{B}|>0$, as long as $|\delta_{B}|=|g\mu_{B}m_{j}B_{z}|$ is much less than the harmonic spacing $h\nu$, or $|B_{z}|\ll|h\nu/g\mu_{B}m_{j}|$.
What may not be obvious is that these now independent conserved quantities act as independent thermodynamic reservoirs in the energy and spin angular momentum degrees of freedom. In Appendices A-E we explicitly show that an ensemble with only spin angular momentum has thermodynamic properties, including: spin analogous heat, work and heat capacity (re-named spin therm, spin labor and waste capacity respectively), unitless temperature $\tau$, and Bose-Einstein and Fermi-Dirac statistics. These properties and more are presented as a summary in Table.2, with them equally applying to any mutually independent conserved quantity.
@@ IV The entropy battery
Bosonic heat in the
thermodynamic limit
@@
Table 2: Summary of spin-analogous thermodynamic properties. In these expressions, $\braket{\bar{J}_{z}}=\braket{J_{z}}/\hbar$ is the unitless average angular momentum, $d_{m}$ is the number of states for the energy, spin $m=E,s$ degrees of freedom respectively, NN is the number of particles, $\varepsilon_{j}=h\nu_{j}$ is the energy of the $j$th vibrational mode with frequency $\nu_{j}$, and $\varepsilon_{D}=h\nu_{D}$ is the Debye energy. Where relevant we have also included references. The waste response is a generalization of the heat capacity to arbitrary conserved quantities, and is coined based on the results of the fluctuation-dissipation theorem, whereby $C_{V}$ and $C_{s}$ are response functions (2). The derivation for each of these expressions for the spin angular momentum conserved quantity is presented in appendices A-E. The first and second laws of generalized thermodynamics apply to closed systems, which is the scenario used in the entropy battery model. The third law describes the generalized thermalization condition of two canonical ensembles.
As these reservoirs are initially independent, the entropy of the ensemble is extensive
The entropy capacity (entropy at infinite unitless temperature $\tau=\infty$) of each reservoir is determined by eqs.(2)-(2), depending on the particles’ statistics. The entropy capacity for a single conserved quantity reservoir is dependent only on the number of microstates $\Omega$, with $\Omega$ dependent on the number of states per particle $d$ and the number of particles $N$.
@@
where $|nd^{\dagger}d\ket{n}\approx$ $\ket{n}$, and $\Omega$ being the Raman Rabi rate, then for the example of a single particle in the initial state
@@
the evolution produces an exchange of heat between the energy and spin degrees of freedom. This exchange liberates heat energy as work by transferring entropy into the spin reservoir, which transforms the spin equivalent of work, spin labor $\mathcal{L}$, into the spin equivalent of heat, spin therm $\mathcal{Q}$. Details of the Raman transition can be found in [12, 13]. Applying this unitary operation to an ensemble of particles, we can begin to realize the extent of possible work extraction. In Fig.1 we show a graphical representation of the entropy transfer process as described above.
Figure 1: Overview of (a) a conventional Carnot heat engine between hot and cold baths, and (b) the entropy battery. In (b), the equivalent conserved quantity ‘heat’ baths with general conserved quantity $\braket{A_{m}}$ are represented by black cubes aligned vertically. Energy couples to spin angular momentum within the entropy battery via a unitary ‘heat engine’, realized experimentally by coherent Raman transitions [12, 13, 14]. Spin labor $\mathcal{L}$ within the entropy battery enables the conversion of heat energy into work. There aren’t any equivalent unitary operations for the other conserved quantities that we are aware of. Entropy from the environmental reservoir is transferred into the battery’s internal reservoirs (energy, spin, etc.), enhancing work extraction until their entropy capacities are saturated or equilibrium is reached. As these other conserved quantity ‘heat’ baths reside within the battery’s particles, the entropy battery achieves a higher per particle work extraction density than possible with conventional heat engines.
@@
We would like to note that the adiabatic Raman transition, which acts as a heat engine, incurs a non-energetic resource cost. Specifically, the engine requires coherence to convert heat into work, as required by the second law of generalized thermodynamics (2). As mentioned previously, this coherence is provided by the spin labor contained within the spin polarized ensemble, $\mathcal{L}$. Therefore, the entropy battery can be viewed as a device that exploits the coherence within general quantum resources, in this case spin angular momentum, to enhance work extraction efficiency.
IV.1 Specifics of the entropy battery